OpenStudy (anonymous):

Someone please help!! Calc 1!! Implicit differentiation xy^1/3+y=10

OpenStudy (amistre64):

what have you tried?

OpenStudy (amistre64):

i see product and power rule ... and implicit tends to mean a chain rule as well

OpenStudy (isaiah.feynman):

@ninab731

OpenStudy (amistre64):

ab + c = k a'b + ab' + c' = k'

OpenStudy (anonymous):

I am trying to get online on my computer am on my iPad !!

OpenStudy (anonymous):

ok am here !! so far i have \[\frac{ d }{ dx }(xy^1/3)+\frac{ d }{ dx }(y)=\frac{ d }{ dx}(10)\]

OpenStudy (amistre64):

im not fond of the d/dx notation for most things

OpenStudy (amistre64):

10'=0 y' = y' work your product rule

OpenStudy (anonymous):

\[1(\sqrt[3]{y})+x(\frac{ 1 }{ 3y ^{\frac{ 2 }{ 3 }} })+y'=0\]

OpenStudy (amistre64):

whenever we implicit a y variable, the chain rule pops out a y'

OpenStudy (amistre64):

\[1(\sqrt[3]{y})+x(\frac{ 1 }{ 3y ^{\frac{ 2 }{ 3 }} })~y'+y'=0\]

OpenStudy (isaiah.feynman):

\[\frac{ x }{ 3\sqrt[3]{y^{2}} }y' +y' = -\sqrt[3]{y}\]

OpenStudy (isaiah.feynman):

You can solve for y' from here.

OpenStudy (anonymous):

thank you both!